# Universality of gravitational radiation from magnetar magnetospheres

### v1

@@ V Implications for detectability
which depends on the inclination from the rotation axis to the line-of-sight ($\iota$), antenna-pattern response functions for the two polarisation states ($F_{+}$ and $F_{\times}$), polarisation angle $\psi$, and phase evolution $\tilde{\phi}(t)$ (see Ref. [52] for details). In expression (12), $S_{n}$ denotes the (instrument-dependent) noise spectral density of the detector.
@@ V.1 Dipole fields
$\approx 10^{22}\left(\frac{Q^{22}}{10^{38}\text{ g cm}^{2}}\right)^{2}\left(\frac{1\text{ s}}{P}\right)^{6}\text{ erg s}^{-1},$
which is always small relative to electromagnetic spindown unless the object has a millisecond spin period, meaning it can be ignored for a slow rotator. For an inclined star with $P\gg 1\text{ ms}$ surrounded by magnetospheric plasma, we thus have [55, 62],
$P\dot{P}\approx\frac{2\pi^{2}B_{p}^{2}R^{6}\left(a+b\sin^{2}\chi\right)}{3c^{3}\left(1-2\mathcal{C}\right)^{4}I_{0}},$ (15)
@@ VI Discussion
In this paper we explore how currents impact on the detectability of GWs from magnetar magnetospheres. Because solutions to equations (1) and (2) can sustain a maximum free energy relative to an untwisted state of only $\approx 30\%$ if there are no magnetic islands [25], GW amplitudes associated with the magnetosphere are confined to a narrow range. The strain (7) is controlled by the combination of $B_{0}^{2}R^{5}$ (expression 10) regardless of the twist particulars (see Fig. 3). It is in this sense we argue for a universal behaviour between objects, akin to others described in the literature for various aspects of the interior [74]. Although detection prospects are marginal for Galactic objects even with the planned, space-based detector DECIGO – most sensitive in the $\lesssim 10$ Hz band – universality is of theoretical interest as probing the near-surface fields of magnetars is difficult owing to timing anomalies, state switching, and other phenomena [29, 30].
With respect to multimessenger connections, suppose that the neutron star undergoes some (sub-)crustal event such that energy is deposited into the magnetosphere. This energy will inject helicity, and thus may lead to a modulation of the characteristic strain until the system relaxes over a diffusion timescale which could reach $\sim 40$ yr for moderate twists in an X-ray dim magnetar with luminosity $\sim 10^{33}\text{ erg s}^{-1}$ (see expression 38 in Ref. [75]). The quadrupole moment may therefore ‘wander’ in a way that can be tied to timing anomalies. Such modulations could be calculated using a set of models, like that we have constructed here, that are sequenced by assuming that the twist decays linearly – as is roughly consistent with post-outburst observations [18, 76, 30]. In principle, signals with sufficient SNR may be disentangled to reveal such features coincident with X-ray activity (though cf. Ref. [77]). Next-generation technology could thus provide detailed information about the ignition mechanism of flares (i.e., whether they are magnetospheric in origin or crustal) and spindown for nearby ($\sim 100$ pc) objects like CHIME J0630.
@@ Acknowledgements
We thank Juan A. Miralles for discussions. We acknowledge support from the Prometeo excellence programme grant CIPROM/2022/13 funded by the Generalitat Valenciana. We also acknowledge support through the grant PID2021-127495NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by the European Union, as well as the Astrophysics and High Energy Physics programme of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1).

### v2

@@ V Implications for detectability
which, as noted in Sec. III.2, depends on the antenna-pattern response functions for the two polarisation states ($F_{+}$ and $F_{\times}$), polarisation angle $\psi$, and phase evolution $\tilde{\phi}(t)$ (see Ref. [52] for details). In expression (14), $S_{n}$ denotes the (instrument-dependent) noise spectral density of the detector. We emphasise again that we discuss only the detectability of the magnetospheric contribution (i.e., setting $h_{0}=h_{0}^{\rm magneto}$), which varies little between models due to the universal behaviour explored throughout. To estimate the detectability of the total system, one should substitute a combined amplitude $h(t)\equiv\sum_{i}\left(F_{+}h^{i}_{+}+F_{\times}h^{i}_{\times}\right)$ [63] into the Fourier integral (14) (see, e.g., Ref. [59] for a discussion).
@@ V.1 Dipole fields
[…]
which is always small relative to electromagnetic spindown unless the object has a millisecond spin period and the fluid quadrupole is very large, meaning it can be ignored for a slow rotator. For an inclined star with $P\gg 1\text{ ms}$ surrounded by magnetospheric plasma, we thus have [57, 67],
$P\dot{P}\approx\frac{2\pi^{2}B_{p}^{2}R^{6}\left(a+b\sin^{2}\chi\right)}{3c^{3}\left(1-2\mathcal{C}\right)^{7/2}I_{0}},$ (17)
@@ VI Discussion
In this paper we explore how currents impact on the detectability of GWs from magnetar magnetospheres. Because solutions to equations (1) and (2) can sustain a maximum free energy relative to an untwisted state of only $\approx 30\%$ if there are no magnetic islands [25], GW amplitudes associated with the magnetosphere are confined to a narrow range. The strain (7) is controlled by the combination of $B_{0}^{2}R^{5}$ (expression 10) regardless of the twist particulars (see Fig. 3). It is in this sense we argue for a universal behaviour between objects, akin to others described in the literature for various aspects of the interior [79]. Although detection prospects are marginal for Galactic objects even with the planned, space-based detector DECIGO – most sensitive in the $\lesssim 10$ Hz band – universality is of theoretical interest as probing the near-surface fields of magnetars is difficult owing to timing anomalies, state switching, and other phenomena [29, 30]. As emphasised in Sec. III.2, the magnetospheric signal sets an effective floor to the GW luminosity and thus our results should be thought of as quantifying detectability in the most conservative scenario. Nevertheless, since the phase portrait is different between the hydromagnetic and magnetospheric pieces (equation 12), the detectability of the latter is largely independent of the former.
With respect to multimessenger connections, suppose that the neutron star undergoes some (sub-)crustal event such that energy is deposited into the magnetosphere. This energy will inject helicity, and thus may lead to a modulation of the characteristic strain until the system relaxes over a diffusion timescale which could reach $\sim 40$ yr for moderate twists in an X-ray dim magnetar with luminosity $\sim 10^{33}\text{ erg s}^{-1}$ (see expression 38 in Ref. [80]). The quadrupole moment may therefore ‘wander’ in a way that can be tied to timing anomalies. Such modulations could be calculated using a set of models, like that we have constructed here, that are sequenced by assuming that the twist decays linearly – as is roughly consistent with post-outburst observations [18, 81, 30]. In principle, signals with sufficient SNR may be disentangled to reveal such features coincident with X-ray activity (though cf. Ref. [82]). Next-generation technology could thus provide detailed information about the ignition mechanism of flares (i.e., whether they are magnetospheric in origin or crustal) and spindown for nearby ($\sim 100$ pc) objects like CHIME J0630 [78].
@@ Acknowledgements
We thank Juan A. Miralles for discussions. AGS acknowledges funding from the European Union’s Horizon MSCA-2022 research and innovation programme “EinsteinWaves” under grant agreement No. 101131233. We acknowledge support from the Prometeo excellence programme grant CIPROM/2022/13 funded by the Generalitat Valenciana. We also acknowledge support through the grant PID2021-127495NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by the European Union, as well as the Astrophysics and High Energy Physics programme of the Generalitat Valenciana ASFAE/2022/026 funded by MCIN and the European Union NextGenerationEU (PRTR-C17.I1).